Consider the inverse demand curve p = 144 - 20 and the cost function C(Q) = 100+ 4Q. If the market were competitive, calculate the incidence of a specific tax, t = 6, that falls on consumers. Calculate the incidence of the same tax if the market were instead a monopoly.

1
Expert's answer
2021-12-08T09:05:25-0500

Competitive market

Demand is P = 144 - 2Q and supply is P = 4

Tax causes supply to shift up and so it becomes P = 10

Before tax quantity is 70 and after tax quantity is 68

Consumers pay 10 and producers receive 4 so entire tax burden is borne by consumers

Welfare is reduced by 0.5*(70 - 68)*6 = $6

Monopoly market

Marginal revenue is MR = 144 - 4Q and MC = 4

Before tax Q = 35 units and P = 74

After tax MC = 10 and so new quantity is 34 units and P = 76

Hence consumers are paying $2 while monopolist is bearing $4 of the tax

Welfare is reduced by 0.5*(70 - 35)*(74 - 4) - 0.5*(68 - 34)*(76 - 10) = 103